1.) 2(y-6)=3y+12-y

2.) 6-2y=7y+13

3.) 3x-5(x-6)=2(10-x)+10
4.) 5n+7=7(n+1)-2n
5.) -4(p+2)+8=2(p-1)-7p+15

Answer by funmath(2933)

1.) 2(y-6)=3y+12-y
2y+2(-6)=3y-y+12
2y-12=2y+12
-2y+2y-12=-2y+2y+12
0-12=0+12
-12=12 No solution, this is inconsistent.
:
2.) 6-2y=7y+13
6-2y+2y=7y+2y+13
6-0=9y+13
6=9y+13
6-13=9y+13-13
-7=9y+0
-7=9y
-7/9=9y/9
-7/9=y
:
3.) 3x-5(x-6)=2(10-x)+10
3x-5(x)-5(-6)=2(10)+2(-x)+10
3x-5x+30=20-2x+10
2x+30=-2x+30
2x+2x+30=-2x+2x+30
4x+30=30
4x+30-30=30-30
4x+0=0
4x=0
4x/4=0/4
x=0
:
4.) 5n+7=7(n+1)-2n
5n+7=7(n)+7(1)-2n
5n+7=7n+7-2n
5n+7=7n-2n+7
5n+7=5n+7 This is an identity, there are infinitly many solutions.
:
5.) -4(p+2)+8=2(p-1)-7p+15
-4(p)-4(2)+8=2(p)+2(-1)-7p+15
-4p-8+8=2p-2-7p+15
-4p+0=2p-7p-2+15
-4p=-5p+13
-4p+5p=-5p+5p+13
p=0+13
p=13
Happy Calculating!!!